solve 8y''-2y'-y=x2+x+1


1
Expert's answer
2021-11-09T13:00:27-0500

Homogeneous differential equation is 8y"-2y'-y=0. -------(1)

Auxiliary equation,

8m2-2m-1=0

m="\frac{1}{2},\frac{-1}{4}"

Therefore, solution of (1),

y(x)="c_1e^{\frac{1}{2}x}+c_2e^{\frac{-1}{4}x}"


Now, finding the particular solution of the given differential equation.

"y_p=\frac{1}{8D^2-2D-1}(x^2+x+1)\\
=(-1)[1+(2D-8D^2)]^{-1}(x^2+x+1)\\
=(-1)[1-(2D-8D^2)+(2D-8D^2)^2-(2D-8D^2)^3-...](x^2+x+1)\\\text{Eliminate higher degree terms}\\
=(-1)[1-(2D-8D^2)+(2D-8D^2)^2](x^2+x+1)\\
=(-1)[x^2+x+1-(2x+1)+24]\\
=-x^2+3x-23\\
\text{Therefore, general solution is }\\
y(x)=c_1e^{\frac{1}{2}x}+c_2e^{\frac{-1}{4}x}-x^2+3x-23"